OpenStudy (anonymous):

What are the x-intercepts for the function f(x)=x^2-8x+15?

OpenStudy (anonymous):

y=15

OpenStudy (anonymous):

15

OpenStudy (solomonzelman):

set f(x)=0 \[x^2-8x+15=0\]\[x^2-8x+16=1\]\[(x-4)^2=1\]continue please....

OpenStudy (anonymous):

it's x intercept not y intercept

OpenStudy (anonymous):

I figured it out, you just do the star method I guess.

OpenStudy (anonymous):

star method? who?

OpenStudy (solomonzelman):

jakejoker, are you talking to me or yourself? Why intercept is when x=0 and x-intercept is when y=0. DUDE !!! wow !

OpenStudy (anonymous):

x-interception is when x=0 and y-interception is when y=0 ???

OpenStudy (anonymous):

OMG. I'm F** up

OpenStudy (anonymous):

sorry dude